Two angles are complementary if the sum of their measures is 90 degrees. Find two complementary angles such that one of the angles is 94 degrees less than 3 times the other angle.
Answer by macston(5194)

.
x=other angle; 3x-94 degrees=one of the angles
x+(3x-94 degrees)=90 degrees
4x-94 degrees=90 degrees
4x=184 degrees
x=46 degrees ANSWER 1: One angle is 46 degrees
3x-94 degrees=3(46 degrees)-94 degrees=138 degrees-94 degrees=44 degrees
ANSWER 2: The other angle is 44 degrees.